Using provided data, solve for present value. When calculating 1/(1+I/Y)^N, round your answer to four decimal places. Use your rounded answer to calculate PV; enter PV rounded to the nearest dollar. When entering (1+I/Y) into the formula, be sure to solve for (1+I/Y) and then enter the result rounded to two decimal places.

Facts

PV:

FV:

$140,000

Annual Interest Rate:

12%

I/Y:

1%

n: (Number of years)

5

N: (Number of compounding periods)

60

Compounded (Y):

Monthly

Formula Method

PV

=

FV

X

1/

(1+I/Y)

^N

=

X

1/

=

X

=
Answer #1

faise Value 28 140000 Anterest = 1% per month No.d lesiod 2 60 brosent Value a fair value (1+x)^N 140000 (17.01) 60 1,40000 (

(1+I/Y)= 1+.01= 1.01 (two decimal)

1/(1+I/Y)^60 = 1.8167 (4 decimal)